# On the origins of counterrotating stellar disks in TNG50. I

### v1

@@ 5.2 Compact ex-situ CRDs
This subtype of CRD is compact in extent but contains a significant fraction of ex-situ stellar particles. By definition, compact ex-situ CRDs have more than 20% of their counterrotating stars accreted from satellite galaxies. We identify 7 galaxies in this category (CR-14, discussed earlier, falls just below the 20% threshold). These are marked with red rhombus symbols in Figures 6 and 7 (see also Appendix Figures 10 and 11).
@@
As shown on Fig. 5, past mergers are of great importance in the assembly of compact ex-situ CRDs. In several cases, a single satellite merger provided the bulk of the accreted stellar particles for the CRD. For example, CR-4, CR-6, CR-7, CR-15, CR-19 and CR-23 each have one dominant progenitor galaxy contributing the majority of the ex-situ component in the CRD. Only one case, CR-2, has two significant contributors to its ex-situ stellar mass. Often, these mergers also enhanced in-situ star formation in the host. As a result, the infalling satellites not only donated stars but also triggered new star formation in retrograde orbits, as evidenced by the star formation histories. Indeed, Fig. 7 shows clear star formation peaks coincident with the merger times for these galaxies. We recall the reader that, in all panels, the most significant contributor is shown with a highlighted arrow.
@@
Figure 8: The orbit of the satellite that is continuing its interaction with CR-0 at z=0. We follow the evolution of the satellite orbit to a lookback time of 6 Gyr. Top panel: evolution of the satellite galactocentric distance from the host. Bottom panel: evolution of the angle between the angular momentum of the disk and the satellite’s orbital angular momentum. Orbits near $180^{\circ}$ are counterrotating, while those near $0^{\circ}$ are co-rotating.
@@
Notably, some of the most massive satellite interactions in our entire sample are associated with compact ex-situ CRDs. CR-2, CR-4, CR-6, and CR-7 experienced some of the highest-mass merger events (in terms of total satellite-to-host mass ratio) of all the galaxies with CRDs (mass ratio ¿ 40%). These massive satellites are strongly affected by dynamical friction, rapidly sinking to the host galactic center. The fact that these major interactions resulted in relatively compact counterrotating disks (confined to the bulge region) indicates that the stars brought in by the merger, and those induced by associated star formation bursts, remained concentrated in the central regions of the galaxy. It is also worth noting that, as in the compact in-situ cases, these CRDs contain an old significant in-situ component tracing back to the formation of the galaxy’s primordial inner regions.
@@
5.3 Extended In-situ CRDs
@@
Extended CRDs are the minority in our TNG50 sample. By definition, these systems have $R_{95}^{\rm CRD}/R_{\rm opt}>0.5$, indicating that their counterrotating stellar disks extend beyond half the optical radius of the host galaxy. Among the 26 identified CRDs, only five meet this criterion. All five of these are predominantly of in-situ origin and are marked with orange circle symbols in Fig. 6 and 7 (see also Fig. 10 and 11).
@@
The extended in-situ CRDs, CR-0, CR-3, CR-5, CR-10, and CR-17, formed the bulk of their counterrotating stars within the main galaxy, with relatively low ex-situ stellar mass fractions. Specifically, CR-0, CR-3, CR-5, and CR-10 each have approximately $11-17\%$ of their CRD stellar mass contributed by satellites, while CR-17 is almost entirely in-situ, with less than $2\%$ ex-situ material. Despite their in-situ dominance, all five systems show clear signs that interactions played a pivotal role in their formation, either by supplying retrograde gas or perturbing the host. As illustrated in Fig. 7, which traces the stellar age distributions and merger events for each CRD, these galaxies experienced one or more significant interactions around the epoch of their main counterrotating star formation. Notably, and in contrast to the majority of CRDs in our sample, the stellar distribution in these five systems is significantly extended.
@@
Examples of this CRD type are shown in Fig. 3. Two peculiar cases are CR-0 and CR-5, which present a torus-like spatial distribution of the CRD component. Both present a concentration $R_{95}^{\rm CRD}/R_{50}^{\rm CRD}\leq 2$ (see Sec. 4). In the CR-0 case, its CRD’s star formation history (Fig. 7) shows at least three distinct peaks. The oldest peak, around 12 Gyr ago (lookback time), corresponds to the very early assembly of the galaxy and is associated with both gas accretion during these early times and a burst associated to a very significant merger, highlighted with a green arrow. The next peak, near 9 Gyr ago, is associated with the infall of a second significant satellite (blue arrow in Fig. 7). This interaction induced a burst star formation, contributing to the CRD in the inner galactic region , as sceen in the previous CRD type. The most recent rise in CR-0’s counterrotating star formation begins $\sim 3$ Gyr ago and continues toward the present day. Interestingly, CR-0 is experiencing an ongoing interaction with a massive satellite that had it first pericentric passage $\approx 4$ to 5 Gyr ago. The satellite has not fully merged at $z=0$ and continues to orbit the host. In Fig. 8, we analyze the orbit of this infalling satellite. The top panel shows the time evolution of the satellite’s galactocentric distance, while the bottom panel the evolution of the angle between the angular momentum of the disk and the satellites’s orbital angular momentum. Orbits near $180^{\circ}$ are counterrotating, those near $90^{\circ}$ are polar, and those near $0^{\circ}$ are co-rotating. The figure clearly shows that the satellite is accreted in a counterotating orbit with respect to the main disc. The satellite is relatively gas rich, with a gas-to-total mass fraction at infall of 17%. The gas stripped on this counterotating orbit fuels the more recent star formation episode of its CRD.
@@
CR-5 also presents a torus-like CRD structure. Its star formation peak occurred around 9.5 Gyr ago and coincides with a major interaction involving a massive gas-rich satellite. This satellite had its first pericentric passage at $\sim 11$ Gyr ago and deposited a significant amount of co-rotating gas at the outermost pre-exising disk region, that subsequently formed the stellar torus-like structure. Interestingly, after the formation of this structure, CR-5 experienced a significant accretion of gas, counterrotating with respect to the pre-existing structure, giving rise to the present-day overall disk. This is shown in the top panel of Fig. 9. Here we show, with blue lines, the time evolution of the disk’s angular momentum vector orientation with respect to its orientation at the present-day, $J_{\rm disk}^{t_{0}}$. In addition, we show with an orange line, the time evolution of the CRD angular momentum vector orientation with respect to $J_{\rm disk}^{t_{0}}$. We can clearly see that the CRD and the overall disk remain well aligned until $t_{L}=7.5$ Gyr. After this time, both component become progressively more misaligned. This is due to the later accretion of misaligned, as shown by the black line which follows the time evolution of the gas angular momentum vector orientation, $J_{\rm gas}$, with respect to $J_{\rm disk}^{t_{0}}$. To compute $J_{\rm gas}$ we considered, at every snapshot, all gas cells enclosed within a sphere of radius equal to the present-day $R_{\rm opt}$. This clearly indicates the significant accretion of misaligned cold gas within $4.5\lesssim t_{L}\lesssim 7.5$ Gyr. Within this period of time the increasing misalignment of the gas component is subsequently followed by what results in the present-day corrotating disk.
@@
CR-10 and CR-17 present a similar formation mechanism of the CRD component. In the case of CR-10, its CRD formed through multiple nearly simultaneous satellite interactions at early time. Four different satellites (2 of them with mass ratios over 25%, other 2 being 12%) interact in a narrow time window with the host. The most massive interacting satellites are indicated by the red arrows in Fig. 7. A central burst of star formation, associated with these interactions, contributed to both the corotating and counterotating components of the final disc. These satellites also contributed with both, stellar particles and extended gas component that gives rise to the present-day CRD. Interesting, as shown in the middle panel of Fig. 9, the present-day CRD arises as a corrotating component. However, the posterior accretion of misaligned gas, starting at $t_{L}\approx 11.5$ Gyr gives rise to a new rotating component that, at the present-day, represents the corrotating disk. CR-17 CRD, instead, almost entirely formed in-situ (only 2% ex-situ) within two early burst of SF; a first episode at a lookback time $t_{L}\approx 10$ Gyr and a second episode 7 Gyr ago. We note that during that period of time, the SF activity of the present-day co-rotating component was almost negligible. Indeed, as shown in Fig. 7, the present-day co-rotating disk starts to grow in mass after $t_{L}\approx 7$ Gyr. In the bottom panel of Fig. 9 we can cleary see that the CRD is born as a corrotating component. However, after $t_{L}\approx 7.5$ Gyr, significant accretion of counterotating gas gives rise the present-day corrotating disc.
@@
Finally, CR-3 is an outlier in several aspects. It’s the closest extended CRD to being ex-situ; about 17% of its stellar mass is directly accreted from satellites. CR-3 also extends unusually far into the outer disk (high $R_{95}^{\rm CRD}/R_{\rm opt}$). Its formation history (Fig. 7) shows a prominent star formation peak around 8 Gyr ago, coinciding with the infall of multiple significant satellites in rapid succession. Two of these satellites were fairly massive ($\approx$ 40% mass ratio). The ex-situ stellar component of this CRD originates from these satellites. The interactions not only deposited stars into the CRD but also funneled gas into the host, which subsequently formed stars on retrograde orbits (hence the substantial in-situ contribution).
@@
Figure 9: Time evolution of the angle between the angular momentum of the disk at z=0 and the disk (blue), CRD (orange) and gas within $R_{\rm opt}$ (black). The angles are obtained using the dot product between the angular momentum vectors as the bottom panel of Fig. 8.
@@
In summary, all four extended in-situ CRDs can trace their formation to interactions: CR-5 and CR-10 each formed during a major burst of multiple simultaneous minor mergers, CR-0’s CRD grew through a combination of early accretion and an ongoing major interaction, and CR-17’s CRD was built via a single low mass merger. However the reasons for counterrotation vary, CR-0 is due to the ongoing accretion of the infalling satellite in a counterrotating orbit. CR-5, CR-10 and CR-17 all have later accretion of misaligned gas that gives rise to a younger stellar disk that ends up dominating over the older CRD. Finally, CR-3 is mainly associated with the simultaneous accretion of significant satellites that contributed with both stars and gas in counterotating orbits. The accretion of counterotating gas that gives rise to some of these CRDs will be studied in detail in a follow-up article.
@@ 6 Discussion
Our study, based on fully cosmological simulations, indicates that counterrotating stellar disks (CRDs) are rare but present in $\lesssim 10\%$ of MW-mass galaxies at $z=0$. And only $\lesssim 1\%$ host an extended CRD. These CRDs contribute only a small fraction of their hosts’ stellar mass (typically $\lesssim 2\%$) and tend to be spatially compact (confined well within the optical radius). A key result is that all but one of these CRDs are dominated by in-situ star formation, with the majority of stars in each counterrotating disk having formed within the galaxy’s potential well rather than being accreted. This is true even in cases with a non-negligible ex-situ stellar contribution. In line with this, most CRDs have distinctly timed star formation episodes in their history – sharp bursts that coincide with some perturbation. This is in agreement with observations, where CRDs are a rare occurrence, encompassing only $\approx 1\%$ of galaxies with ionized gas in MaNGA (Bao et al. 2022, Bevacqua et al. 2022). Interestingly, we find that the simulated CRDs are older, on average, than their present-day corotating stellar disks. This is in contrast with observational studies, where they find that CRDs are mostly younger than the main stellar disk (Pizzella et al. 2014, Gasymov et al. 2025). We also find that, typically, galaxy models with CRDs have the gas and the younger stellar disk component (typically dominant) co-rotating with respect to each other. This in agreement with observations of CRDs in late-type galaxies (Bao et al. 2022, Bao et al. 2024, Gasymov et al. 2025).
@@
Our analysis reveals multiple formation channels for CRDs with varying degrees of external influence. The simulated galaxies exhibit a spectrum of CRD types, from significant merger-driven structures to more internal early star forming episodes. For example, one notable CRD involves an extended retrograde disk (CR-3) built primarily by the accretion of multiple satellite galaxies in quick succession. This event contributed with both stellar particles and a significant amount of gas mass in retrograde orbits, producing one of the most extended CRDs in the sample. There are other four extended in-situ CRDs which, despite having nearly all their counterrotating stars formed in-situ, can each be traced to at least one significant encounter that induced their formation. One of these cases is currently forming stars in its CRD as a result of the prolonged accretion of a satellite on a retrograde orbit. In contrast, the other three, older CRDs are remnants of a previous accretion of misaligned gas that gave rise to the currently dominant younger co-rotating stellar disk. Episodes of gas misalignment giving rise to a younger stellar disk has been observed in past studies (Pizzella et al. 2004; Vergani et al. 2007; Gasymov et al. 2025). Similarly, we identify compact ex-situ CRDs (eight cases); i.e., compact and retrograde disk components with a significant ex-situ contribution. In general, these ex-situ star particles are associated to one significant progenitor. The associated interactions often enhanced in-situ star formation in the inner regions of the host (Barnes & Hernquist 1996, Tissera et al. 2002, Rupke et al. 2010, Bustamante et al. 2018), contributing to both the overall and the CR stellar distribution. Finally, the most common variety in TNG50 are compact in-situ CRDs. These 14 cases are diverse in origin: roughly half formed at very early times, coeval with the galaxy’s initial bulge formation and without any clear external trigger. A few others in this category show evidence of accretion events that induced bursts of in-situ star formation within the central region, that contributed to both the co-rotating and counterrotating components of the final stellar disk.
@@
Despite the variety of formation channels, the in-situ stars’ dominance indicates that gas with a misaligned (retrograde) angular momentum was accreted into the galaxy’s disk along its history. In some cases, this can be directly associated with interactions. Indeed, for the merger-induced CRDs, infalling satellites delivered not just stars but also fresh gas that was subsequently funneled into the host on retrograde orbits. In other cases, we find that misaligned gas accretion takes place after the present-day CRD formation, resulting in a younger stellar disk that ends up dominating over the older CRD. The mechanism behind such accretion will be further studied in a follow-up work.
@@
In summary, nearly all (all but one) simulated CRDs are dominated by stars formed in-situ, even in systems shaped by mergers. Importantly, ‘in-situ’ here includes stars formed within the host from gas stripped or supplied by companions. Interactions often play a role in triggering or shaping the CRD. However, they are not strictly required in all cases. A small fraction of CRDs, especially compact ones, appear to arise naturally during the early phases of galaxy formation, without any recent accretion event. Thus, our results indicate that, although most counterrotating stellar disks are dominated by in-situ star formation, their emergence is frequently linked to external processes. Interactions and the accretion of misaligned gas are common triggers that supply the conditions for retrograde star formation.

### v2

@@ 5.2 Compact ex situ CRDs
This subtype of CRD is compact in extent but contains a significant fraction of ex situ stellar particles. By definition, compact ex situ CRDs have more than 20% of their counterrotating stars accreted from satellite galaxies. We identified seven galaxies in this category (CR-14, discussed earlier, falls just below the 20% threshold). These galaxies are marked with red rhombus symbols in Figures 8 and 9 (see also Appendix Figures 12 and 13).
@@
As shown on Fig. 7, past mergers are of great importance in the assembly of compact ex situ CRDs. In several cases, a single satellite merger provided the bulk of the accreted stellar particles for the CRD. For example, CR-4, CR-6, CR-7, CR-15, CR-19, and CR-23 each have one dominant progenitor galaxy contributing the majority of the ex situ component in the CRD. Only one case, CR-2, has two significant contributors to its ex situ stellar mass. Often, these mergers also enhanced in situ star formation in the host. As a result, the infalling satellites not only donated stars but also triggered new star formation in retrograde orbits, as evidenced by the SFHs. Indeed, Fig. 9 shows clear star formation peaks coincident with the merger times for these galaxies. We recall that, in all panels, the most significant contributor is shown with a highlighted arrow.
@@
Figure 10: Orbit of the satellite that continues its interaction with CR-0 at z=0. We followed the evolution of the satellite orbit to a lookback time of 6 Gyr. Top panel: Evolution of the satellite galactocentric distance from the host. Bottom panel: Evolution of the angle between the angular momentum of the disk and the satellite’s orbital angular momentum. Orbits near $180^{\circ}$ are counterrotating, while those near $0^{\circ}$ are corotating.
@@
Notably, some of the most massive satellite interactions in our entire sample are associated with compact ex situ CRDs. In particular, CR-2, CR-4, CR-6, and CR-7 experienced some of the largest mass merger events (in terms of total satellite-to-host mass ratio) of all the galaxies with CRDs (mass ratio ¿ 40%). These massive satellites are strongly affected by dynamical friction, and they rapidly sink to the host galactic center. The fact that these major interactions resulted in relatively compact coplanar CRDs (confined to the bulge region) indicates that the stars brought in by the merger and those induced by associated star-formation bursts remained concentrated in the central regions of the galaxy. It is also worth noting that, as in the compact in situ cases, these CRDs contain an old and significant in situ component tracing back to the formation of the galaxy’s primordial inner regions.
@@
5.3 Extended in situ CRDs
@@
Extended CRDs are the minority in our TNG50 sample. By definition, these systems have $R_{95}^{\rm CRD}/R_{\rm opt}>0.5$, indicating that their counterrotating stellar disks extend beyond half the optical radius of the host galaxy. Among the 26 identified CRDs, only five meet this criterion. All five of these are predominantly of in situ origin and are marked with orange circle symbols in Fig. 8 and 9 (see also Fig. 12 and 13).
@@
The extended in situ CRDs, CR-0, CR-3, CR-5, CR-10, and CR-17, formed the bulk of their counterrotating stars within the main galaxy, with relatively low ex situ stellar mass fractions. Specifically, approximately $11-17\%$ of the CRD stellar mass of CR-0, CR-3, CR-5, and CR-10 was contributed by satellites, while CR-17 is almost entirely in situ, with less than $2\%$ of ex situ material. Despite their in situ dominance, all five systems show clear signs that interactions played a pivotal role in their formation, either by supplying retrograde gas or perturbing the host. As illustrated in Fig. 9, which traces the stellar age distributions and merger events for each CRD, these galaxies experienced one or more significant interactions around the epoch of their main counterrotating star formation. Notably, and in contrast to the majority of CRDs in our sample, the stellar distribution in these five systems is significantly extended.
@@
We show examples of this CRD type in Fig. 5. Two peculiar cases are CR-0 and CR-5, which present a torus-like spatial distribution of the CRD component. Both present a concentration $R_{95}^{\rm CRD}/R_{50}^{\rm CRD}\leq 2$ (see Sec. 4). In the CR-0 case, the SFH of its CRD (Fig. 9) shows at least three distinct peaks. The oldest peak, around 12 Gyr ago (lookback time), corresponds to the very early assembly of the galaxy and is associated with both gas accretion during these early times and a burst associated with a very significant merger, which we highlight with a green arrow in the figure. The next peak, near 9 Gyr ago, is associated with the infall of a second significant satellite (blue arrow in Fig. 9). This interaction induced a burst of star formation that contributed to the CRD in the inner galactic region, as seen in the previous CRD type. The most recent rise in CR-0’s counterrotating star formation began $\sim 3$ Gyr ago and continues to the present day. Interestingly, CR-0 is experiencing an ongoing interaction with a massive satellite that had its first pericentric passage $\approx 4$ to 5 Gyr ago. The satellite has not fully merged at $z=0$ and continues to orbit the host. In Fig. 10, we analyze the orbit of this infalling satellite. The top panel shows the time evolution of the satellite’s galactocentric distance, while the bottom panel shows the evolution of the angle between the angular momentum of the disk and the satellites’s orbital angular momentum. Orbits near $180^{\circ}$ are counterrotating, those near $90^{\circ}$ are polar, and those near $0^{\circ}$ are corotating. The figure clearly shows that the satellite is accreted in a counterotating orbit with respect to the main disk. The satellite is relatively gas rich, with a gas-to-total mass fraction at infall of 17%. The gas stripped on this counterotating orbit fuels the more recent star formation episode of the CRD in the host galaxy.
@@
CR-5 also presents a torus-like CRD structure. Its star formation peak occurred around 9.5 Gyr ago and coincides with a major interaction involving a massive gas-rich satellite. This satellite had its first pericenter passage $\sim 11$ Gyr ago, and it deposited a significant amount of corotating gas at the outermost preexisting disk region, which subsequently formed the stellar torus-like structure. Interestingly, after the formation of this structure, CR-5 experienced a significant accretion of gas, counterrotating with respect to the preexisting structure, giving rise to the present-day disk. The rise of the younger dominant stellar disk is shown in the top panel of Fig. 11. In the figure, with blue lines we show the time evolution of the disk’s angular momentum vector orientation with respect to its orientation at the present-day, $J_{\rm disk}^{t_{0}}$. In addition, with an orange line we show the time evolution of the CRD angular momentum vector orientation with respect to $J_{\rm disk}^{t_{0}}$. One can clearly see that the CRD and the overall disk remain well aligned until $t_{L}=7.5$ Gyr. After this time, both components become progressively more misaligned. This is due to the later accretion of misaligned gas, as shown by the black line, which follows the time evolution of the gas angular momentum vector orientation, $J_{\rm gas}$, with respect to $J_{\rm disk}^{t_{0}}$. To compute $J_{\rm gas}$, we considered at every snapshot all gas cells enclosed within a sphere of radius equal to the present-day $R_{\rm opt}$. This clearly indicates the significant accretion of misaligned cold gas within $4.5\lesssim t_{L}\lesssim 7.5$ Gyr. Within this period of time, the increasing misalignment of the gas component is subsequently followed by what results in the present-day corotating disk.
@@
Notably, CR-10 and CR-17 present a similar formation mechanism of the CRD component. In the case of CR-10, its CRD formed through multiple nearly simultaneous satellite interactions at an early time. Four different satellites (two with mass ratios over 25%, and the other two with ratios of 12%) interacted in a narrow time window with the host. We indicate the most massive interacting satellites with red arrows in Fig. 9. A central burst of star formation associated with these interactions contributed to both the corotating and counterotating components of the final disk. These satellites also contributed stellar particles and an extended gas component that gave rise to the present-day CRD. Interestingly, as shown in the middle panel of Fig. 11, the present-day CRD arises as a corotating component. However, the posterior accretion of misaligned gas starting at $t_{L}\approx 11.5$ Gyr gives rise to a new rotating component that, at the present day, represents the corotating disk. In contrast, CR-17 CRD formed almost entirely in situ (only 2% ex situ) within two early bursts of star formation: a first episode at a lookback time of $t_{L}\approx 10$ Gyr and a second episode 7 Gyr ago. We note that during both episodes of formation in the CRD, the star-formation activity of the present-day corotating component was almost negligible. Indeed, as shown in Fig. 9, the present-day corotating disk starts to grow in mass after $t_{L}\approx 7$ Gyr. In the bottom panel of Fig. 11, it can be clearly seen that the CRD is born as a corotating component. However, after $t_{L}\approx 7.5$ Gyr, significant accretion of counterotating gas gives rise to the present-day corotating disk.
@@
Finally, CR-3 is an outlier in several aspects. It is the closest extended CRD to being ex situ; about 17% of its stellar mass is directly accreted from satellites. CR-3 also extends unusually far into the outer disk (high $R_{95}^{\rm CRD}/R_{\rm opt}$). Its formation history (Fig. 9) shows a prominent star formation peak around 8 Gyr ago, coinciding with the infall of multiple significant satellites in rapid succession. Two of these satellites were fairly massive ($\approx$ 40% mass ratio). The ex situ stellar component of this CRD originates from these satellites. The interactions not only deposited stars into the CRD but also funneled gas into the host, which subsequently formed stars on retrograde orbits (hence the substantial in situ contribution).
@@
Figure 11: Time evolution of the angle between the angular momentum of the disk at z=0 and the disk (blue), CRD (orange), and gas within $R_{\rm opt}$ (black). Angles near $180^{\circ}$ are counterrotating, while those near $0^{\circ}$ are corotating in comparison to the present-day disk. Separations in angle between components represent a misalignment between them.
@@
In summary, all four extended in situ CRDs can trace their formation to interactions: CR-5 and CR-10 each formed during a major burst of multiple simultaneous minor mergers; the CRD of CR-0 grew through a combination of early accretion and an ongoing major interaction; and the CRD of CR-17 was built via a single low mass merger. However, the reasons for counterrotation vary. For CR-0, it is due to the ongoing accretion of the infalling satellite in a counterrotating orbit, whereas CR-5, CR-10, and CR-17 all have later accretion of misaligned gas that gives rise to a younger stellar disk that ends up dominating over the older CRD. Finally, CR-3 is mainly associated with the simultaneous accretion of significant satellites that contributed both stars and gas in counterotating orbits. The accretion of counterotating gas that gives rise to some of these CRDs will be studied in detail in a follow-up article.
@@ 6 Discussion
Our study, based on fully cosmological simulations, indicates that CRDs are rare but present in $\lesssim 10\%$ of MW-mass galaxies at $z=0$, and that only $\lesssim 1\%$ host an extended CRD. These CRDs comprise only a small fraction of their hosts’ stellar mass (typically $\lesssim 2\%$) and tend to be spatially compact (confined well within the optical radius). This is opposed to lenticular and early-type disk galaxies where the counter-rotating components found in them tend to be of a higher mass fraction and are also more common in simulations (Khoperskov et al. 2021, Lu et al. 2021) and in observations (Gasymov et al., 2025), where coplanar as well as non-coplanar counterrotating disks are considered. A key result of our work is that all but one of the CRDs we find are dominated by in situ star formation, with the majority of stars in each CRD having formed within the galaxy’s potential well rather than being accreted. This is true even in cases with a non-negligible ex situ stellar contribution. In line with this, most CRDs have distinctly timed star formation episodes in their history – sharp bursts that coincide with some perturbation. This is in agreement with observations, where counterrotating stellar disks are a rare occurrence, encompassing only $\approx 1\%$ of galaxies with ionized gas in MaNGA (Bao et al. 2022, Bevacqua et al. 2022). Interestingly, we find that the simulated CRDs are on average older than their present-day corotating stellar disks. This is in contrast to observational studies, where counterrotating stellar disks are mostly found to be younger than the main stellar disk (Pizzella et al. 2014, Gasymov et al. 2025). We also find that galaxy models with CRDs typically have the gas and the younger stellar disk component (typically dominant) corotating with respect to each other. This in agreement with observations of counterrotating stellar disks in late-type galaxies (Bao et al. 2022, Bao et al. 2024, Gasymov et al. 2025).
@@
Our analysis revealed multiple formation channels for CRDs with varying degrees of external influence. The simulated galaxies exhibit a spectrum of CRD types, from significant merger-driven structures to more internal early star forming episodes. For example, one notable CRD involves an extended retrograde disk (CR-3) built primarily by the accretion of multiple satellite galaxies in quick succession. This event contributed both stellar particles and a significant amount of gas mass in retrograde orbits, producing one of the most extended CRDs in the sample. There are four other extended in situ CRDs that despite having nearly all their counterrotating stars formed in situ, can each be traced to at least one significant encounter that induced their formation. One of these cases is currently forming stars in its CRD as a result of the prolonged accretion of a satellite on a retrograde orbit. In contrast, the other three older CRDs are remnants of a previous accretion of misaligned gas that gave rise to the currently dominant younger corotating stellar disk. Episodes of gas misalignment giving rise to a younger stellar disk have been observed in past studies (Pizzella et al. 2004; Vergani et al. 2007; Gasymov et al. 2025). Similarly, we identified compact ex situ CRDs (eight cases), i.e., compact and retrograde disk components with a significant ex situ contribution. In general, these ex situ star particles are associated with one significant progenitor. The associated interactions often enhanced in situ star formation in the inner regions of the host (Barnes and Hernquist 1996, Tissera et al. 2002, Rupke et al. 2010, Bustamante et al. 2018), contributing to both the overall and the CR stellar distribution. Finally, the most common variety of CRDs in TNG50 are compact in situ CRDs. We identified 14 cases, and they are diverse in origin: Roughly half formed at very early times, coeval with the galaxy’s initial bulge formation and without any clear external trigger. A few others in this category show evidence of accretion events that induced bursts of in situ star formation within the central region that contributed to both the corotating and counterrotating components of the final stellar disk.
@@
Despite the variety of formation channels, the dominance of the in situ stars indicates that gas with a misaligned (retrograde) angular momentum was accreted into the galaxy’s disk along its history. In some cases, this can be directly associated with interactions. Indeed, for the merger-induced CRDs, infalling satellites delivered not just stars but also fresh gas that was subsequently funneled into the host on retrograde orbits. In other cases, we found that misaligned gas accretion takes place after the present-day CRD formation, resulting in a younger stellar disk that ends up dominating over the older CRD. The mechanism behind such accretion will be further studied in a follow-up work.
@@
In summary, nearly all (all but one) simulated CRDs are dominated by stars formed in situ, even in systems shaped by mergers. Importantly, “in situ” here includes stars formed within the host from gas stripped or supplied by companions. Interactions often play a role in triggering or shaping the CRD. However, they are not strictly required in all cases. A small fraction of CRDs, especially compact ones, appear to arise naturally during the early phases of galaxy formation and show no evidence of any recent accretion event. Thus, our results indicate that although most counterrotating stellar disks are dominated by in situ star formation, their emergence is frequently linked to external processes. Interactions and the accretion of misaligned gas are common triggers that supply the conditions for retrograde star formation.